OpenStudy (anonymous):

The sum of two nonnegative numbers is 20. Find the numbers if the sum of their squares is as large as possible; as small as possible.

OpenStudy (anonymous):

as large as possible you do in your head since you cannot tell the two apart, largest is when they are equal (by symmetry)

OpenStudy (anonymous):

you can use algebra if you like i suppose make your teacher happy maybe find the minimum of \(x^2+(20-x)^2\)

OpenStudy (anonymous):

man i got that backwards min you do in your head not max

OpenStudy (anonymous):

i apologize, but i forgot how to do maxes and mins....like, how do you find the different combinations?

OpenStudy (anonymous):

for min they are both 10, for max one is 20, the other is 0

OpenStudy (anonymous):

one number is \(x\) say, then since the total is 20, the other must be \(20-x\) sum of the squares is \[x^2+(20-x)^2=2x^2-40x+400\] parabola that face up

OpenStudy (anonymous):

minimum is at the vertex first coordinate of the vertex is \(-\frac{b}{2a}=-\frac{-40}{4}=10\)

OpenStudy (anonymous):

max is at the endpoint of the domain domain is \([0,20]\) and this is entirely symmetric about 10, so max is where one is 20 and the other is zero

OpenStudy (anonymous):

Oh ok, thanks so much!

OpenStudy (anonymous):

yw